Question 4.2 points. The data below are presented for a fictional economy. Using the base year as 2008 and assuming that a market basket consists of 4 oranges, 2 lemons, and 1 grapefruit, I have entered some information for you to aid your calculations. 2008 0.10 0.20 2009 0.12 $0.21 $0.50 Price of orange S0.13 Price of lemon $0.50 $0.21 $0.54 Price of grapefruit S0.54 Cost of market basket S0.10 4+ $0.20 2+ $0.50*1-$1.30 ($1.30/S1.30)x100 - 100 CPI A) Calculate the CPI in 2009, 2010, and 2011. Show your work B) Using the CPI numbers you calculated in part A, calculate the inflation rate in 2009, 2010, and 2011.
Answer #1

(A) Cost of market basket in 2009 = $(0.12)(4)+ $(0.21)(2) + $(0.50)(1) = $(0.48 + 0.42 +0.50)= $ 1.40

CPI in 2009 = (cost of market basket in 2009/ cost of market basket in base year) 100= $(1.40/1.30)100= 107.69

Cost of market basket in 2010 = $(0.12)(4) +$(0.21)(2) + $(0.54)(1) = $(0.48 + 0.42 + 0.54) =$1.44

CPI in 2010 = (Cost of market basket in 2010/ cost of market basket in base year) 100 = $(1.44/1.30)100= 110.77

Cost of market basket in 2011 = $(0.13)(4) +$(0.21)(2) + $(0.54)(1) = $(0.52 +0.42 +0.54) = $1.48

CPI in 2011 = (cost of market basket in 2011/ cost of market basket in base year) 100= $(1.48/1.30)100 = 113.85

(B) Inflation rate in 2009 = [(CPI in 2009 -CPI in 2008)/ CPI in 2008 ] 100 = [(107.69-100/ 100] 100= 7.69%

Inflation rate in 2010 = [(CPI in 2010- CPI in 2009) /CPI in 2009] 100 = [(110.77-107.69)/ 107.69 ] 100 = 2.86%

Inflation rate in 2011 =[(CPI in 2011- CPI in 2010)/ CPI in 2010] 100 = [(113.85- 110.77)/ 110.77] 100= 2.78%